# Topology-induced symmetry breaking demonstrated in antiferromagnetic magnons on a Möbius strip

### v4

@@ III Analysis and Discussion
which is apparently different from that under the TBC. These two subspaces, in spite of their partial overlap, are distinct and cannot be transformed into each other by a linear combination of the coefficients.
One may wonder what happens when the system size approaches infinity. In this limit, the spectral shift of the $\alpha$ branch vanishes so the two branches become degenerate in the eigenvalues. Nevertheless, the eigenvectors (hence the eigenspace) for each allowed $k$ remain structurally distinct from their counterparts associated with the PBC. In other words, the eigenspaces under TBC and PBC do not merge into each other as the system approaches infinity. In particular, the $\pi$ phase shift of the $\alpha$ mode at $k=0$ is topological protected and independent of the system size.
Now let us inspect the manifestation of the proposed TISB. It is established that in easy-axis AFM materials such as MnF2 [19, 32], $R$ and $L$ are degenerate in energy in the absence of magnetic fields. In easy-plane AFM materials such as NiO, the existence of hard-axis anisotropy explicitly breaks the rotational symmetry in the spin Hamiltonian and lifts the degeneracy, rendering $L-R$ and $L+R$ the magnon eigenmodes [20, 21, 22]. In these common scenarios, the magnon eigenstates exhibit the same symmetry with the spin Hamiltonian, as we only consider the excited states while excluding the ground state that is subject to the spontaneous symmetry breaking and recasts the broken symmetry in the Goldstone modes. By contrast, the Hamiltonian (5) we adopted preserves the local rotational symmetry (with respect to the local easy axis), which is explicitly broken in the magnon eigenstates ascribing to the Möbius TBC. This intriguing phenomenon is a direct manifestation of the TISB, i.e., non-trivial topology in the real space alone leads to lifted degeneracy in the eigenvalues, as well as reduced symmeetry in the eigenspace of elementary excitations, without the aid of symmetry-breaking interactions. Because a quantum of angular momentum associated with the $R$-($L$-)circular mode is $+\hbar$ ($-\hbar$), the TISB we found should be followed by the suppression of longitudinal magnon spin currents on a Möbius strip.
Figure 4: (a) Illustration of the standing wave configurations for the lowest three $\alpha$ ($\beta$) modes with an odd (even) number of nodes $n$. The vertical axis indicates the amplitude of spin oscillation, where a sign difference corresponds to a $\pi$ phase difference. (b) Illustration of the elliptical spin trajectories (on site $i=1$) that smoothly deform with a varying $J_{AF}$. Other sites (not shown) follow a similar pattern of change.
@@ IV Exotic Phase
What the $\alpha$ and $\beta$ branches of modes share in common is that the elliptical orbits of spin precessions are always elongated in the $x$-direction (longitudinal direction of the strip). This important property, according to Eq. (18), is attributed to $q_{\alpha(\beta)}^{1}>q_{\alpha(\beta)}^{2}$, where $q_{\alpha(\beta)}^{1}=Q_{\alpha(\beta)}+J_{AF}/2$ and $q_{\alpha(\beta)}^{2}=Q_{\alpha(\beta)}-J_{AF}/2$ with $J_{AF}>0$. On this point, it is instrumental to notice that if $J_{AF}<0$, i.e., the AFM inter-chain exchange coupling turns ferromagnetic, then the opposite relation will be true: $q_{\alpha(\beta)}^{1}<q_{\alpha(\beta)}^{2}$, leading to elongated trajectories in the $y$-direction (transverse direction of the strip). However, an immediate problem with the $J_{AF}<0$ regime is that the AFM ground state may become unstable, which will be reflected as the magnon spectra (eigenvalues) touching zero.
One may naively expect a phase transition when $J_{AF}$ flips sign such that the AFM state yields to the ferromagnetic state and forms a domain wall [23, 24]. However, a closer inspection reveals a quite counter-intuitive result. According to Eqs. (15) and (16), we have
@@
and similarly, the minimum of $\omega_{\beta}$ occurs at $k=0$, where $\min[\omega_{\beta}]=\min[\omega_{\alpha}]$. One can tell from Eq. (26) that the AFM ground is preserved even for $J_{AF}<0$, as long as $K+J_{AF}>0$. A phase transition must take place at $J_{AF}=-K$, where magnons become infinitely soft and proliferate, melting the AFM configuration.
The surprising region of $-K<J_{AF}<0$ features an exotic phase in which the AFM ground state can stand even though the inter-chain and intra-chain exchange interactions are both ferromagnetic. As schematically illustrated in Fig. 4(b), the elliptical trajectories of sublattice spins deform smoothly while $J_{AF}$ varies from positive to negative, until the threshold $J_{AF}=-K$. An ideal circular polarization for each sublattice is reached at $J_{AF}=0$. In the exotic phase with $-K<J_{AF}<0$, the major axes of spin precessions indeed lie in the $y$-direction. Nonetheless, the magnon eigenspace is still a disparate subspace comparing to that dictated by the PBC, regardless of the sign of $J_{AF}$. While the linear polarization status of the Néel vector and the total spin vector remain the same for both $J_{AF}>0$ and $-K<J_{AF}<0$, the magnitudes of their oscillations are noticeably different in the two phases, which can be pictorially inferred from Fig. 4(b) and be rigorously quantified via the same set of formulas in Sec. II and III.
@@ V Outlook
Even though we have demonstrated the TISB in the AFM magnons on a Möbius strip, the mechanism itself is general and can manifest in other contexts with different quasiparticles or TBCs. For example, we anticipate that the eigenmodes of phonon excitations on a Möbius strip to be linearly polarized while the circularly-polarized chiral phonons are suppressed by the Möbius topology. For photons that are governed by Maxwell’s equations, the axis of circular polarization is parallel to the momentum $\bm{k}$, which defines a distinct geometry compared to ours, calling for a separate investigation. For tight-binding electrons on a Möbius strip, while the Hamiltonian is very similar to ours, the fermionic statistics brings a fundamental distinction in the diagonalization [which does not involve $\sigma_{z}$ as that in Eq. (15)]. Under alternative TBCs such as a Klein bottle, even the AFM magnons could acquire new features of TISB beyond what we showed in this work. Therefore, our findings could greatly inspire a broader research endeavor in the near future highlighting the profound impact of real-space topology on the physical nature of elementary excitations.

### v5

@@ III Analysis and Discussion
where $0<\eta_{\rm small}<\eta_{\rm large}$ and the minus sign indicates a $\pi$ phase difference for $\bm{S}_{A}$ and $\bm{S}_{B}$ shown in Fig. 3(d). This is apparently different from that under the TBC. These two subspaces, in spite of their partial overlap for some special parameters, are generally distinct and cannot be transformed into each other by a linear combination of the coefficients, given that $q_{\alpha}^{1(2)}\neq q_{\beta}^{1(2)}$. Starting from a 4D Hilbert space, one needs to specify the PBC ($A\rightarrow A$ and $B\rightarrow B$) or TBC ($A\rightarrow B$ and $B\rightarrow A$) when reconciling site $i$ with $i+N$, thereby projecting the original 4D space into a chosen 2D subspace. In Sec.IV, one can further appreciate the distinctions between the two different 2D subspaces.
One may wonder what happens when the system size approaches infinity. In this limit, the spectral shift of the $\alpha$ branch vanishes so the two branches become degenerate in terms of eigenvalues. Nevertheless, the eigenvectors (hence the eigenspace) for each allowed $k$ remain structurally distinct from their counterparts associated with the PBC. In other words, the eigenspace conforming with TBC does not coalesce with the eigenspace under PBC as the system approaches infinity. In particular, the accumulative $\pi$ phase shift of the $\alpha$ mode is topological protected and independent of the system size.
Figure 4: (a) Illustration of the standing wave configurations for the lowest three $\alpha$ ($\beta$) modes with an odd (even) number of nodes $n$. The vertical axis indicates the amplitude of spin oscillation, where a sign difference corresponds to a $\pi$ phase difference. (b) Illustration of the elliptical spin trajectories (on site $i=1$) that smoothly deform when $J_{AF}$ varies. Other sites (not shown) follow a similar pattern of change.
Now let us inspect the manifestation of the proposed TISB. It is established that in easy-axis AFM materials such as MnF2 [19, 32], $R$ and $L$ are degenerate in energy in the absence of magnetic fields. In easy-plane AFM materials such as NiO, the existence of hard-axis anisotropy explicitly breaks the rotational symmetry in the spin Hamiltonian and lifts the degeneracy, rendering $L-R$ and $L+R$ the magnon eigenmodes [20, 21, 22]. In these common scenarios, the magnon eigenstates exhibit the same symmetry with the spin Hamiltonian, as we only consider the excited states while excluding the ground state that is subject to the spontaneous symmetry breaking and recasts the broken symmetry in the Goldstone modes. By contrast, the Hamiltonian (5) we adopted preserves the local rotational symmetry (with respect to the local easy axis), which is explicitly broken in the magnon eigenstates ascribing to the Möbius TBC. This intriguing phenomenon is a direct manifestation of the TISB, i.e., non-trivial topology in the real space alone leads to lifted degeneracy in the eigenvalues, as well as reduced symmetry in the eigenspace of elementary excitations, without the aid of symmetry-breaking interactions. Because a quantum of angular momentum associated with the $R$-($L$-)circular mode is $+\hbar$ ($-\hbar$), the TISB we found should be followed by the suppression of longitudinal magnon spin currents on a Möbius strip.
@@ IV Exotic Phase
While the major axis of the elliptical orbits for $\alpha$ modes lie in the $x$-direction (longitudinal direction of the strip), they are in the $y$-direction for the $\beta$ modes. This important property, according to Eq. (18), is attributed to $q_{\alpha(\beta)}^{1}>q_{\alpha(\beta)}^{2}$, where $q_{\alpha(\beta)}^{1}=Q_{\alpha(\beta)}+J_{AF}/2$ and $q_{\alpha(\beta)}^{2}=Q_{\alpha(\beta)}-J_{AF}/2$ with $J_{AF}>0$. On this point, it is instrumental to notice that if $J_{AF}<0$, i.e., the AFM inter-chain exchange coupling turns ferromagnetic, then the opposite relation will be true: $q_{\alpha(\beta)}^{1}<q_{\alpha(\beta)}^{2}$, leading to elongated trajectories in the $y$-direction ($x$-direction) for the $\alpha$ ($\beta$) modes. However, an immediate problem with the $J_{AF}<0$ regime is that the AFM ground state may become unstable, which will be reflected as the magnon spectra (eigenvalues) touching zero.
One may naively expect a phase transition when $J_{AF}$ flips sign such that the AFM state yields to the ferromagnetic state and necessarily forms a domain wall [23, 24]. However, a careful inspection reveals something counter-intuitive. According to Eqs. (15) and (16),
@@
and similarly, the minimum of $\omega_{\beta}$ occurs at $k=0$, where $\min[\omega_{\beta}]=\min[\omega_{\alpha}]$. One can tell from Eq. (26) that the AFM ground is preserved even for $-K<J_{AF}<0$. A phase transition must take place at $J_{AF}=-K$, where magnons become infinitely soft and proliferate, melting the AFM configuration and driving the ground state into a ferromagnetic domain wall (note that a uniform ferromagnetic configuration is forbidden by the Möbius boundary condition).
The surprising region of $-K<J_{AF}<0$ features an exotic phase in which the AFM ground state can stand even though the inter-chain and intra-chain exchange interactions are both ferromagnetic. As schematically illustrated in Fig. 4(b), the elliptical trajectories of sublattice spins deform smoothly while $J_{AF}$ varies from positive to negative, until the threshold $J_{AF}=-K$. An ideal circular polarization for each sublattice is reached at $J_{AF}=0$. In the exotic regime $-K<J_{AF}<0$, the principal axes of spin precession for both the $\alpha$ and $\beta$ modes indeed swap relative to the $J_{AF}>0$ case. Nonetheless, the magnon eigenspace is still a disparate subspace comparing to that dictated by the PBC, regardless of the sign of $J_{AF}$.
The $\alpha$ and $\beta$ modes in this exotic phase has a unique feature which is fundamentally impossible for the conventional linearly-polarized modes $L+R$ and $L-R$: The trajectories of $\bm{S}_{A}$ and $\bm{S}_{B}$, when being projected onto the local $x-y$ plane, meet along the major axis and become opposite to each other along the minor axis. While the linear polarization status of the Néel vector and the total spin vector remain the same for both $J_{AF}>0$ and $-K<J_{AF}<0$, the magnitudes of their oscillations are noticeably different in the two phases, which can be pictorially inferred from Fig. 4(b) and be rigorously quantified via the same set of formulas in Sec. II and III.
@@ V Outlook
Even though we have demonstrated the TISB in the AFM magnons on a Möbius strip, the mechanism itself is general and can manifest in other contexts with different quasiparticles or TBCs. For example, we anticipate that the eigenmodes of phonon excitations on a Möbius strip to be linearly polarized while the circularly-polarized chiral phonons are suppressed by the Möbius topology. For photons that are governed by Maxwell’s equations, the axis of circular polarization is parallel to the momentum $\bm{k}$, which defines a distinct geometry compared to ours, calling for a separate investigation. For tight-binding electrons on a Möbius strip, while the Hamiltonian is very similar to ours, the fermionic statistics brings a fundamental distinction in the diagonalization [which does not involve $\sigma_{z}$ as that in obtaining Eq. (15)]. Under alternative TBCs such as a Klein bottle, even the AFM magnons could acquire new features of TISB beyond what we have shown in this work. Therefore, our findings could greatly inspire a broader research endeavor in the near future highlighting the profound impact of real-space topology on the physical nature of elementary excitations.